let me see if I have this correct?
7x-5+3x=6+x-10
7x+3x-5=x+6-10
10x-5=x-4
(10x-5)+(5-x)=(x-4)+(5-x)
(10x-5)+(-x+5)=(-x-4)+(-x+5)
So would -4 be the solution to the equation?
Thank you, Barb Neely

Answer by edjones(8007)

7x-5+3x=6+x-10
7x+3x-5=x+6-10
10x-5=x-4
(10x-5)+(5-x)=(x-4)+(5-x) This is very good. You have correctly added 5-x to both sides. The parentheses are not necessary since you are adding and the signs wont change.
10x-5+5-x=x-4+5-x is better. Now the 5's disappear on the left and the x's disappear on the right.
(10x-5)+(-x+5)=(-x-4)+(-x+5) This is bad. Why did you add 5-x to each side? So you could get all the x's on one side of the equal sign and all the digits on the other. Tidying up on the right side caused you to make an error. (-x-4) should be (x-4). As you know, in addition or subtraction the order of the numbers doesn't have any affect on the answer so reordering them only gives you another chance to make a mistake. It's like rearranging the deck chairs on the Titanic.
.
Here's how I did it.
7x-5+3x=6+x-10
10x-5=x-4
10x-x-5+5=x-x-4+5
9x=1
9x/9=1/9
x=1/9
.
You are improving, Barb.
You'll get it.
.
Ed